An oil drilling company ventures into various locations, and their success or failure is
independent from one location to another. Suppose the probability of a success at any
specific location is 0.25.
i) What is the probability that a driller drills 10 locations and finds 1 success?
ii) The driller feels that he will go bankrupt if he drills 10 times before the first
success occurs. What are the driller’s prospects for bankruptcy?

Answer by stanbon(75887)

Suppose the probability of a success at any
specific location is 0.25.
i) What is the probability that a driller drills 10 locations and finds 1 success?
P(one success in 10) = (10C1)(0.25)^1(0.75)^9 = 0.1877...
--------------------
ii) The driller feels that he will go bankrupt if he drills 10 times before the first success occurs. What are the driller’s prospects for bankruptcy?
----------------
P(no success in 10 trials) = 0.75^10 = 0.0563...
Chances are 563 in 10000 that he will go bankrupt.
===============
Cheers,
Stan H.